Janet is mixing a 15% glucose solution with a 35% glucose soution. This mixture produces 35 liters of a 19% glucose solution. How many liters of the 35% solution did Janet use?
Answer by ewatrrr(24785)


Hi,
mixing a 15% glucose solution with a 35% glucose soution
Let x represent the 35% solution (assigning x to more concentrated solution)
then (35-x) will be the amount of the 15% soluton
.35x + .15(35L-x) = .19*35L
solving for x
.35x + .15*35L -.15x = .19*35L
.20x = .04*35L
x = .15*35L/.20 = 7L, the amount of the 35% solution
checking our answer
28L of the 15% solution (35L - 7L)
.35*7 + .15*28 = 6.65 = .19*35